### v1

@@ Appendix B Suggested student projects
The thermodynamic application of Section 3 suggests several other investigations for a student researcher. Including elastic hard-sphere collisions would affect individual trajectories, but the state variables of the system would not be affected. This is similar to the model of the ideal gas. This might lead to thermalization on a shorter time scale. The container wall can be moved more quickly, allowing work to be done on the system. It would be interesting to explore possible hysteresis effects of a container moving inward and then allowed to expand.
@@
Additionally, these systems can be simulated in situations that include outside forces. For example, linear drag could be applied to one or more of the particles, allowing students to illustrate known results, such as the ultimate dissipation of all energy in a system experiencing linear drag, and the return of the affected particle to its initial position [28]. One could also simulate an $N$-body attractive system released in a uniform gravitational field and allowed to bounce on a hard surface, tracking the motion of the center of mass and the characteristic size of the system as it experiences these outside forces.
A selection of glowscript programs is included [29], which simulate interactions with an external environment. Due to external interactions, these simulations rely on traditional numerical integration, rather than the analytic form of the solutions. As anticipated, these simulation run times scale as $N^{2}$, in contrast with the simulations that make use of the analytic solution, which scale linearly with $N$.

### v3

@@ Appendix B Student Projects
The thermodynamic applications suggest several other investigations for a student researcher. Including elastic hard-sphere collisions, as in an ideal gas model, would affect individual trajectories but the state variables of the system would be unaffected. It might however lead to thermalization on a much shorter time scale.
In this work, particles collide with the a stationary container wall. Collisions with a moving container would model work done on the system. It would be interesting to explore possible hysteresis effects of a container moving inward and then allowed to expand.
One could also investigate the effects of changing the effective temperature of a fixed container. This could be accomplished by making the collisions with the wall inelastic, simulating a point of impact wall velocity. The statistical speed distribution of the effective wall velocity would provide the temperature of the wall.
@@
Additionally, these systems can be simulated in situations that include outside forces. For example, linear drag could be applied to one or more of the particles, allowing students to illustrate known results. One could also simulate an $N$-body attractive system released in a uniform gravitational field and allowed to bounce on a hard surface, tracking the motion of the center of mass and the characteristic size of the system as it experiences these outside forces. Selected python programs demonstrating these external interactions have been prepared using glowscriptbartz_trinkets . Due to external interactions, these simulations rely on traditional numerical integration, rather than the analytic form of the solutions. As anticipated, these simulation run times scale as $N^{2}$, in contrast with the simulations that make use of the analytic solution, which scale linearly with $N$.
Appendix C Analysis of log-log plots of thermodynamic quantities
In this supplemental material, we present log-log plots of all thermodynamic quantities from the main text, for both attractive and repulsive cases. A straight line fit on a log-log plot suggests a power-law relationship, with the slope indicating the exponent.
While the best-fit lines appear visually convincing, residual analysis reveals significant issues. In key cases, residuals exhibit systematic behavior, contradicting the assumption of a linear log-log fit. This is evident in the attractive case for pressure vs.volume (Fig.7) and the repulsive case for temperature vs.volume (Fig.9).
In the attractive case, temperature data fluctuations dominate the residuals (Figs. 6, 8). These fluctuations are temporally correlated, as seen in the wave-like residual patterns, reflecting non-random, systematic deviations. Conversely, pressure data shows less fluctuation and more systematic residual behavior.
In the repulsive case, pressure data shows greater fluctuation than temperature (Figs. 10, 11). While the pressure is autocorrelated in time, residuals do not exhibit clear systematic deviation or patterns. This inconsistency underscores why the relationships among thermodynamic variables require further analysis.
Figure 6: Attractive case: Log(temperature) vs. Log(volume) and residuals. The log-log plot shows temperature as a function of volume, with the best-fit line indicating an apparent power-law relationship. However, the residuals exhibit a wave-like pattern, indicating systematic deviations and suggesting that the power-law fit is not adequate.
Figure 7: Attractive case: Log(pressure) vs. Log(volume) and residuals. The log-log plot suggests a power-law relationship, but the residuals show clear systematic trends, indicating that the linear fit on the log-log plot does not accurately describe the data.
Figure 8: Attractive case: Log(pressure) vs. Log(temperature) and residuals. However, the residuals exhibit a wave-like pattern, indicating systematic deviations and suggesting that the power-law fit is not adequate.
Figure 9: Repulsive case: Log(temperature) vs. Log(volume) and residuals. The log-log plot indicates an apparent power-law relationship, but the systematic trend in the residuals undermines the validity of the fit.
Figure 10: Repulsive case: Log(pressure) vs. Log(volume) and residuals.
Figure 11: Repulsive case: Log(pressure) vs. Log(temperature) and residuals.